The car costs $12590 and is 20% off. What was the original price?
answer $15737.50
no comprendo. please explain!

Answer by JimboP1977(311)

Ok do you understand what 20% off something means?
It means that 20% of that price is taken away from its original price.
So we are saying that the original price minus 20% of its original price equals its new price.
A good rule to remember is to find a percentage of something: find 1% first and then multiply by the percentage you want to find.
For example if you want to find 20% of the original price you would divide the original price by 100 to get 1% and then multliply by 20 to get 20%
Mathematically:
x = original price
y = new price
x-(20/100)*x = y
So in the question we know y but do not know x. So we need to rearrange to find x.
Take the common factor out of x
x(1-(20/100)) = y
x(4/5) = y
x = y/(4/5)
x = 12590/(4/5)
x = 15737.5
Another way to think about is is that the original price multiplied by 80% (100%-20%) equals the reduced price?
Does that help? Or is it more confusing? Please get back to me.